Fermi National Laboratory
 

Accelerator Update for July 18, 2007 - July 20, 2007

 

Wednesday July 18

The day shift began with Operations monitoring the Tevatron (TeV) store 5559 (75.0E30) and the Recycler stash (101.0E10), with the Antiproton Source (Pbar) stacking (50.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams, and with MiniBooNE and the Meson MTest experiments taking beam. NuMI is off until after the summer shutdown.

At 8:45 AM, a Booster RF station (BRF2) tripped off. It reset.

At 10:46 AM, a Linac RF station (LRF2) tripped off. It reset.

MI experts began a study period at 12:09 PM. A Linac expert began tuning the Linac. Pbar experts conducted a reverse proton study during an extended transfer of antiprotons.

At 1:07 PM, BRF13 & 14 tripped off when a vacuum pump (IPRF23) tripped off. Everything reset okay.

The MI studies ended at 2:12 PM.

At 2:31 PM, LRF2 tripped off. It reset.

Operations reported at 3:58 PM that the Recycler had stashed antiprotons three times during the day shift.

At 5:18 PM, MIRF12 tripped off. It reset.

At 8:38 PM, BRF2 tripped off. It reset.

At 9:04 PM, Operations reported that after the last stash, the Debuncher momentum band would not turn on. Also, the Pbar Accumulator emittances began to grow. Operators contacted experts.

At 9:11 PM, a TeV sector D49 separator sparked. There was no significant loss of luminosity.

At 9:22 PM, a Tevatron Electron Lens (TEL1) tripped off. It reset.

MTest finished taking beam for the night at 9:29 PM.

Operations terminated store 5559 at 10:04 PM.

TeV experts began a study period at 10:20 PM.

Operations reported at 10:22 PM that stacking had resumed. An expert reported that the Pbar Debuncher and Accumulator problems were due to a PLC that had locked up. It was reset and everything returned to normal.

At 11:14 PM, Operations reported a power glitch due to a thunderstorm. Numerous devices in Linac, MI, Recycler, and the TeV tripped off.

At 11:51 PM, Operations reported that they couldn’t get LRF3 to turn on. An operator contacted experts for help.

Operations reported at midnight that the Recycler had stashed twice during the evening shift.

Thursday July 19

The midnight shift began with Operations monitoring the stash (249.7E10) and the stack (28.4E10), and with all beam off due to a problem with LRF3.

TeV studiers gave up for the night. Operations put the TeV into a dry squeeze at 12:12 AM.

Operations reported at 1:05 AM that Linac experts had LRF3 back up and running. Pbar and MiniBooNE resumed taking beam.

At 1:39 AM, Operations reported that a TeV sector A4 vacuum pump had tripped off. Operations prepared the TeV for an access.

Vacuum technicians reported at 5:30 AM that they had replaced two TeV vacuum pumps, one at E4 and the other at A4.

After interlock experts repaired a gate interface card for Meson M01, Operations established beam to MTest at 7:49 AM.

Operations reported at 8:03 AM that the Recycler had stashed three times during the midnight shift.

Operations put the TeV through another dry squeeze at 8:43 AM.

RF experts reported at 8:46 AM that BRF2 was down due to a bad Power Amplifier (PA).

Operations put the TeV into shot setup at 9:23 PM.

At 9:28 AM, Operations reported that they had lost the MI Power Supply interlock Loop, which in turn opened the VCBs. Operations contacted an expert who had MI back on by 9:45 AM.

Pbar resumed stacking at 9:46 AM.

After experts resolved a problem with an energy mismatch between MI and the Recycler, and a Recycler kicker problem, Operations established store 5562 at 12:03 PM with an initial luminosity of 196.16E30.

A MI expert began a one-hour study period at 1:52 PM.

The Neutron Therapy Facility (NTF) reported at 2:45 PM that they had treated their last patient for the day.

At 3:05 PM, LRF3 tripped off. It reset.

Operations reported at 3:57 PM that the Recycler had stashed twice during the day shift.

At 10:25 PM, a Booster vacuum pump (IPRF23) tripped off and caused BRF13 & 14 to trip off. Everything reset.

Friday July 20

The midnight shift began with Operations monitoring store 5562 (50.25E30) and the stash (184.8E10), with Pbar stacking (40.7E10), and with NuMI and MiniBooNE taking beam.

At 12:38 AM, LRF3 tripped off for the first of three times in 2.5 hours. It eventually reset each time.

At 6:44 AM, a Booster vacuum pump (IPRF17) tripped off and caused BRF7 & 8 to trip off. Everything reset.

Operations terminated store 5562 at 8:01 AM.

Operations reported at 8:01 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Friday and the Weekend

The plans for today are to repair the E4 wet engine. The Run Coordinator will allow a few short jobs in the TeV tunnels and studies in Pbar and MI during the wet engine repair. After the wet engine work is completed the plans for the rest of the day and the weekend are to stack, stash, and store, but he will allow TeV studies between stores.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/23/2007    email Fermilab